OpenStudy (anonymous):

Find the volume of a culinder with a radius of 2 meters and a height of 4 meters

OpenStudy (diyadiya):

\[\text{Volume of a cylinder }= pi \times (radius)^2 \times height \]

OpenStudy (diyadiya):

Here radius =2m

OpenStudy (diyadiya):

Height =4m

OpenStudy (campbell_st):

Use the formula r = radius, h = height \[V = pi r^2h\]

OpenStudy (diyadiya):

substitute it to the formula \[Volume = \pi *(2)^2 *4\]

OpenStudy (anonymous):

well i did 3.14* 4*4 ?

OpenStudy (diyadiya):

You're Right :)

OpenStudy (anonymous):

ok thanks i thought i did it wrong

OpenStudy (diyadiya):

Nope thats right :)